Breanna is 6 years older than Sidney. In two years Breanna will be twice as old as Sidney. HOw old is each?
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

B = S+6
B+2 = 2(S+2)
B+2 = 2S+4
S+6+2 = 2S+4
S+8 = 2S+4
S = 4
Sidney is 4 and Breanna is 10
:
Happy learning,
John

Answer by MathTherapy(10557)

Breanna is 6 years older than Sidney. In two years Breanna will be twice as old as Sidney. HOw old is each?
Correct answer: 
Any other answer is INCORRECT, and should be ignored!